A commercial ship is heading due North



at 12.0 km/hr but drifts westward at 5.0



km/hr. determine the magnitude and



direction of the resultant velocity of the



ship.

1
Expert's answer
2021-12-24T11:32:03-0500

We can find the magnitude of the resultant velocity from the Pythagorean theorem:


"v_{res}=\sqrt{v_x^2+v_y^2}=\sqrt{(5\ \dfrac{km}{h})^2+(12\ \dfrac{km}{h})^2}=13\ \dfrac{km}{h}."

We can find the direction of the resultant velocity from the geometry:


"\theta=tan^{-1}(\dfrac{v_y}{v_x})=tan^{-1}(\dfrac{12\ \dfrac{km}{h}}{5\ \dfrac{km}{h}})=67.4^{\circ}\ N\ of\ W."